TB 13-52 Which of the following measures would assist... Which of the following measures would assist in assessing the profitTB 13-71 Judging only from the ratios given, which of... Judging only from the ratios given, which of the following clothingTB 13-79 If a companys If a companys P/E ratio is 24 and the companys EPS is $1.50 then the companys share price is: poinNet sales revenue Expenses Interest Income tax expense Net cash from operations Assets end of current year Liabilities end ofTB 13-88 Calculate points Net sales revenue Cost of goods sold Average accounts receivable Average inventory Average plant, p
Answer #1

11. Current ratio is a Liquidity ratio, Debt-to-asset ratio is a Solvency ratio and Receivables turnover ratio is a Efficiency ratio. Whereas Fixed asset turnover ratio measures profitability of the company. Hence the answer is Fixed asset turnover ratio.

12. Receivables turnover ratio is a ratio that measures how efficiently the accounts receivables are collected. Inventory turnover ratio is a ratio that measures how many times a company has sold its inventory. For a company to have least cash flow problems, the receivables turnover ratio and the inventory turnover ratio should be high. Hence the answer is Company C, who has a receivables turnover ratio of 10 and a inventory turnover ratio of 10.

13. P/E Ratio - 24 times

Earnings per share (EPS) - $1.50

Company's Share Price = P/E Ratio * EPS = 24 * $1.50 = $36.00

14. Current Market Price = $15 per share

Earnings per share = Net Income/Common Stock outstanding = ($900,000 - $500,000 - $10,000 - $90,000)/40,000 = $7.5

Price to Earnings Ratio = Current Market Price/Earnings per share = $15/$7.5 = 2.00 times

15. Turnover = $345,000

Average accounts receivable = $32,500

Accounts Receivable turnover ratio = Turnover/Average accounts receivable = $345,000/$32,500 = 10.62 times